Barton Industries expects next year's annual dividend, D1, to be $2.30 and it expects dividends to grow at a constant rate g = 4.8%. The firm's current common stock price, P0, is $22.80. If it needs to issue new common stock, the firm will encounter a 5.1% flotation cost, F. Assume that the cost of equity calculated without the flotation adjustment is 12% and the cost of old common equity is 11.5%. What is the flotation cost adjustment that must be added to its cost of retained earnings
1
Expert's answer
2019-09-12T11:05:18-0400

The flotation cost adjustment is: D1/(P0×(1 - F)) + g = 2.3/(22.8×(1 - 0.051)) + 0.048 = 0.1542 or 15.42%.